A sztochasztikus folyamatok gyakorlaton való részvétel

dátum:  II. 12. II. 19. II. 26. III. 4. III. 11. III. 18. III. 25.

diákok:

Aradi Bernadette  +  +  +  +  +  +  −
Barkován Anett  +  +  +  +  +  +  −
Barna Timea  +  −  +  +  +  +  +
Bodnár Nelli  +  +  +  +  +  +  +
Csubák Balázs  −  +  −  −  −  −  −
Dobány Máté  +  +  −  +  +  +  −
Fista Norbert  −  +  +  +  −  +  +
Kész Timea  +  +  +  +  +  +  −
Kontor Ibolya  +  +  +  +  +  +  +
Kovács Viktor  +  +  +  +  −  −  +
Mályi Nikoletta  +  −  +  +  +  +  −
Mezei Anita  +  +  +  +  −  +  +
Nagy Beáta  +  +  +  −  +  +  −
Sebestyény Csilla  +  +  +  −  +  +  −
Sikolya Kinga  +  +  +  +  +  +  +
Soróczky Mihály  +  +  +  +  +  +  +
Szabó Angéla  +  +  +  +  +  +  −
Szokol Patricia  +  +  +  +  −  +  +
Tóth Veronika  +  +  +  −  +  +  −
Tóth Zoltán  +  +  +  +  +  +  −
Vas Renáta  +  +  +  +  +  +  +
Weinémer Andrea  +  +  +  +  +  +  +

dátum:  IV. 1. IV. 8. IV. 15. IV. 22. IV. 29. V. 13.

diákok:

Aradi Bernadette  +  −  +  +  +  +
Barkován Anett  +  +  +  +  −  −
Barna Timea  −  +  +  +  +  −
Bodnár Nelli  +  +  +  +  +  +
Csubák Balázs  −  −  −  −  −  −
Dobány Máté  +  +  −  +  +  +
Fista Norbert  −  +  +  +  +  +
Kész Timea  +  +  +  +  +  +
Kontor Ibolya  +  +  +  +  +  −
Kovács Viktor  +  +  +  −  +  −
Mályi Nikoletta  +  +  −  +  +  +
Mezei Anita  −  +  +  +  +  +
Mizsei Zsolt  −  +  +  +  +  +
Nagy Beáta  +  +  +  +  −  +
Sebestyény Csilla  +  +  +  +  −  +
Sikolya Kinga  +  +  +  +  +  +
Soróczky Mihály  +  +  +  +  +  +
Szabó Angéla  +  +  +  +  +  +
Szokol Patricia  +  +  +  +  +  +
Tóth Veronika  +  +  +  +  −  +
Tóth Zoltán  −  +  +  +  +  +
Vas Renáta  +  +  −  +  +  +
Weinémer Andrea  +  +  +  +  +  +